OpenStudy (anonymous):

Add. Simplify by collecting like radical terms if possible, assuming that all expressions under radicals represent nonnegative numbers. sqrt10a +2sqrt90a^3 I'm confused. How do I go about solving this?

OpenStudy (anonymous):

the square roots go all the way across the problem

OpenStudy (anonymous):

is the answer 10?

jimthompson5910 (jim_thompson5910):

First, we can simplify \[\sqrt{90a^3}\] like so \[\sqrt{90a^3}\] \[\sqrt{9*10a^2*a}\] \[\sqrt{9}*\sqrt{10}*\sqrt{a^2}*\sqrt{a}\] \[3*\sqrt{10}*a*\sqrt{a}\] \[3a*\sqrt{10a}\] So \[\sqrt{90a^3}=3a*\sqrt{10a}\] ------------------------------- Now this means that the original expression \[\sqrt{10a}+2\sqrt{90a^3}\] turns into this \[\sqrt{10a}+2*3a*\sqrt{10a}\] \[\sqrt{10a}+2*3a*\sqrt{10a}\] \[\sqrt{10a}+6a\sqrt{10a}\] \[(1+6a)\sqrt{10a}\] So in the end, \[\sqrt{10a}+2\sqrt{90a^3}=(1+6a)\sqrt{10a}\]